Solve the following equation.

dx/y-xz=dy/yz+x=dz/x^2+y^2


1
Expert's answer
2021-06-15T04:36:53-0400

"\frac{dx}{y-xz}=\frac{dy}{yz+x}=\frac{dz}{x^2+y^2}"


"\frac{xdx}{x(y-xz)}=\frac{ydy}{y(yz+x)}=\frac{dz}{x^2+y^2}"


"\frac{xdx-ydy}{-z(x^2+y^2)}=\frac{dz}{x^2+y^2}"


"x^2-y^2+z^2=c_1"


"\frac{ydx+xdy}{x^2+y^2}=\frac{dz}{x^2+y^2}"


"xy-z=c_2"


General solution:

"F(x^2-y^2+z^2,xy-z)=0"